OpenStudy (anonymous):

Is 4x – 3 = 19 true, false, or open?

OpenStudy (anonymous):

Well, it depends on what the value of x is. Read the equation like this - four times some number, minus three equals nineteen. So for this to be true, all values for x have to work. But lets say x equals 1, then 4(1)-3 is 1, and 1 does not equal nineteen, so it cannot be true. However, for it to be false, NO values for x will work to make the equation true. But, notice, if you plug (22/4) as x, then it will simplify to 19=19, which is true. So, this equation will be true sometimes, and false other times.

OpenStudy (hba):

4x-3=19 4x=22 x=22/4 x=11/2

OpenStudy (hba):

Now Put In Question

OpenStudy (hba):

4x-3=19 4(11/2)-3=19 22-3=19 19=19 L.H.S=R.H.S True